OpenStudy (anonymous):

Find the equation of the line that is a perpendicular bisector of the segment with the endpoints of (-3,6) and (11,-2). So, in word document answer the problem by setting up a statement and by setting up a statement and reason table. Please Help

OpenStudy (anonymous):

1) find the midpoint of the line segment 2) find the slope of the line

OpenStudy (anonymous):

you got the midpoint?

OpenStudy (anonymous):

yeah (8,4)?

OpenStudy (anonymous):

hmmmm no you added them up, but you forgot to divide by 2

OpenStudy (anonymous):

my bad (4,2) then right

OpenStudy (anonymous):

yup now how about the slope?

OpenStudy (anonymous):

need help with that one, or is it ok?

OpenStudy (anonymous):

its ok I almost got it

OpenStudy (anonymous):

let me know what you get there are two more steps after that

OpenStudy (anonymous):

i gave you medal cuz you are so cool

OpenStudy (anonymous):

\[\frac{ -8 }{ 14 }\]

OpenStudy (anonymous):

yes, better known as \(-\frac{4}{7}\)

OpenStudy (anonymous):

next step, find the slope of the perpendicular line do you know it?

OpenStudy (anonymous):

perpendicular line means it intersects at 90 degrees correct?

OpenStudy (anonymous):

yes, but more to the point you need to know what the slope of a line perpendicular to the line with slope \(-\frac{4}{7}\) is

OpenStudy (anonymous):

that would make it \[\frac{ 4 }{ 7 }\] right or nah?

OpenStudy (anonymous):

nah flip it

OpenStudy (anonymous):

ok \[\frac{ 7 }{ 4}\] ?

OpenStudy (anonymous):

yes one more step the slope is \(\frac{7}{4}\) and the point is \((4,2)\) use the "point slope" formula to find the equation of the line

OpenStudy (anonymous):

you know it?

OpenStudy (anonymous):

ok so \[y-2=\frac{ 7 }{ 4 } (x-4)?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then multiply out etc

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

ok so \[y=\frac{ 7 }{ 4 }x-5\]

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

Awesome thank you so much

OpenStudy (anonymous):

yw you did all the work